OpenStudy (anonymous):

X^3-8x^2+19x-12 C = 3 is a zero what are the remaining zeros ? What is this in factored form ?

OpenStudy (anonymous):

divide by x - 3 this gives x^2 - 5x + 4 = (x-4)(x -1) zeros are 3 , 4 an 1 f(x) = (x-3)(x-4)(x-1)

OpenStudy (nenadmatematika):